# UAV-Enhanced Combination to Application: Comprehensive Analysis and Benchmarking of a Human Detection Dataset for Disaster Scenarios

### v1

@@ Abstract
Unmanned aerial vehicles (UAVs) have revolutionized search and rescue (SAR) operations, but the lack of specialized human detection datasets for training machine learning models poses a significant challenge. To address this gap, this paper introduces the Combination to Application (C2A) dataset, synthesized by overlaying human poses onto UAV-captured disaster scenes. Through extensive experimentation with state-of-the-art detection models, we demonstrate that models fine-tuned on the C2A dataset exhibit substantial performance improvements compared to those pre-trained on generic aerial datasets. Furthermore, we highlight the importance of combining the C2A dataset with general human datasets to achieve optimal performance and generalization across various scenarios. This points out the crucial need for a tailored dataset to enhance the effectiveness of SAR operations. Our contributions also include developing dataset creation pipeline and integrating diverse human poses and disaster scenes information to assess the severity of disaster scenarios. Our findings advocate for future developments, to ensure that SAR operations benefit from the most realistic and effective AI-assisted interventions possible.
@@ 1 Introduction
The advancement of UAVs, colloquially known as drones, has signaled a new era in the field of emergency response and disaster management. With their unparalleled agility and ability to provide an aerial perspective, drones have rapidly become indispensable assets in the arsenal of SAR operations worldwide. These technological marvels significantly improve the efficiency and effectiveness of missions aimed at locating and aiding people in disaster-hit areas [kucukayan2024yolo]. Drones can have a significant impact on minimizing the aftermath of disasters through time efficiency, making a difference in survival and fatality rates.
@@
In critical scenarios, the lack of datasets with the necessary detail for training advanced machine learning algorithms hinders the optimization of drone capabilities. The primary focus of current resources is on general object detection, which does not provide the depth required to tackle the unique challenges presented by disaster-affected areas [wang2024lovenas]. This limitation significantly restricts the development and refinement of deep learning algorithms capable of accurately identifying human presences amidst the diverse obstacles presented by rubble, floodwater, or fire-damaged territories.
Recognizing the critical need for a specialized dataset, our research introduces a novel dataset explicitly designed to improve human detection capabilities in disaster scenarios through drone vision. This dataset is created by combining human posture images with disaster scene backgrounds, resulting in an intricate collection of images that simulate the diverse settings faced during real-world SAR operations. By integrating human figures into a variety of disaster scenes, our approach aims to mirror the array of challenges that SAR drones are likely to face, facilitating the development of more robust and effective machine learning models tailored for disaster response applications [nunez2024transformer].
Our endeavor is inspired by the growing field of disaster management technology, where recent breakthroughs highlight the revolutionary potential of machine learning and computer vision in improving SAR operations [Aryal]. Our research focuses on creating a challenging dataset to train AI models for detecting partially occluded humans, a common scenario in disaster-stricken environments. Such occlusions, resulting from individuals being trapped under debris or obscured by various objects, represent a significant hurdle for human detection systems. Existing datasets scarcely address this challenge, primarily due to the inherent difficulty in replicating these complex scenarios accurately. Our dataset, therefore, signifies a pioneering step towards addressing this gap, offering a resource that simulates the conditions under which SAR operations unfold in reality.
We aim to enhance the precision and reliability of drone-operated human detection in emergency scenarios, underscoring the nuanced requirements of effective disaster response strategies. The construction of this dataset involved a process of incorporating elements of obstruction to emulate the visibility challenges frequently encountered in real-world disaster settings. This approach facilitates the training of AI models that are not only adept at recognizing human forms in clear view, but are also capable of inferring the presence of individuals in less than optimal visibility conditions. Such abilities are necessary for quickly finding emergency areas and finding survivors who need help right away, which makes SAR missions much more effective [alsamhi2022uav, hu2022seeing].
@@ 2 Literature Review
Robust datasets are fundamental to developing and training precise machine learning models. The importance of datasets tailored to specific application domains is well established in the literature. Studies such as those by [alali2024techniques] and [tang2024review] emphasize the need for datasets that encompass the intricacies of various disaster scenarios. These resources are pivotal for calibrating UAV-operated detection systems to recognize human subjects under a multitude of conditions. Yet, there remains a scarcity of datasets that accurately mirror the complexities of disaster-hit environments.
Debris and destruction frequently obscure human subjects in disaster zones, biasing UAV detection. According to the literature, existing datasets and models are progressing, but they fall short of providing the granular detail required for reliable detection in such complex circumstances [munawar2021uavs, malandrino2019planning]. A significant body of research, including work by [kang2023residential], emphasizes the pressing need for advancing UAV technology to navigate these obstacles adeptly. Yet, the development of datasets that reflect the reality of partial occlusions in disaster contexts is still in its early stages, indicating a pivotal area for future research.
The evolution of machine learning and computer vision has been enhancing to UAV capabilities. These advancements have paved the way for more nuanced data analysis, crucial for discerning human presence within complex terrains. Even with these improvements, the research shows that algorithms and models still need to be improved, especially to make human detection more reliable [khial4725375online].
One notable effort in this direction is the Search and Rescue Drone (SARD) dataset [sambolek2021automatic], which focuses on human detection in search and rescue operations using drone imagery. The SARD dataset includes images of people in various poses simulating exhausted or injured individuals, captured in non-urban environments. While it provides a valuable resource for developing detection models for SAR scenarios, the SARD dataset does not explicitly include disaster scenes, which pose additional challenges such as debris, occlusions, and clutter. Moreover, although the dataset contains diverse human poses, this information is not directly incorporated into the annotations, limiting its utility for pose-aware detection.
@@
Feature C2A (proposed) LoveNAS [wang2024lovenas] Smoke Scene [wu2023dataset] M4SFWD [wang2024m4sfwd] SARD [sambolek2021automatic]
@@ 3.1 Data Sources and Composition
AIDER (Aerial Image Dataset for Emergency Response Applications): The AIDER dataset [kyrkou2019deep] serves as the foundation for the disaster scene backgrounds. It comprises images from four major disaster types: Fire/Smoke (320 images), Flood (370 images), Collapsed Building/Rubble (320 images), and Traffic Accidents (335 images). These authentic disaster images offer a realistic portrayal of the chaotic and unpredictable conditions typical in emergency scenarios. We did not utilize the 1,200 normal case images to keep the focus on emergency situations. This dataset offers a glimpse into the chaotic and unpredictable environments that characterize disaster scenes, making it an ideal choice for our purposes.
LSP/MPII-MPHB (Multiple Poses Human Body): For the human subjects, we sourced images from the LSP/MPII-MPHB dataset [andriluka20142d, johnson2010clustered], which contains 26,675 images featuring 29,732 instances of human bodies in various poses. This dataset is specifically designed to capture a wide range of human body positions, including bent, kneeling, sitting, upright, and lying, providing the necessary diversity to train models for detecting humans under different conditions. The detailed annotations of human poses in this dataset are critical for training models to recognize human figures in complex disaster environments.
@@ 3.2 Pipeline Steps
Using the U2Net segmentation model [qin2020u2], we isolated human figures from the LSP/MPII-MPHB dataset by removing the background. The U2-Net, short for "U-Squared Net," is a deep neural network known for its powerful performance in salient object detection and image segmentation tasks. It employs a nested U-structure that enhances the learning of local and global features within images, enabling precise segmentation of objects, including human figures, from their backgrounds. This process involved saving each figure with its respective pose in a separate folder, ensuring that the focus remained on the human subject without any background distractions.
@@ 4 Properties of C2A Dataset
The C2A (Combination to Application) dataset is a curated collection specifically designed for advancing human detection disaster scenarios by combining AIDER dataset images (disaster scene backgrounds) and diverse human poses from the LSP/MPII-MPHB dataset. Some of the samples of the dataset are presented in Fig. 1. Comparison of various datasets is shown on Table 2. In this section, we present a comprehensive analysis of the dataset’s properties.
Table 2: Comparison of various datasets [xia2018dota] including the proposed C2A dataset
@@ 4.2 Objects Size
Table 3: Comparison of instance size (in terms of width) distribution of some datasets in aerial images and natural images. Some statistics collected from [xia2018dota]
@@ 5.2 Training Options
The evaluation of the models on the C2A dataset was conducted using NVIDIA A100 GPUs, with a uniform batch size of 24 and an image resolution of 640x640 pixels across 50 epochs. The ADAM optimizer was chosen for its efficiency in handling large datasets and complex image structures. Basic data augmentation techniques, such as flipping and resizing, were employed to enhance model robustness and prevent overfitting. The experiments were facilitated by popular deep learning frameworks, specifically mmDetection [chen2019mmdetection], Detectron2 [wu2019detectron2], and Ultralytics [jocher2022ultralytics], known for their high performance in object detection tasks. These frameworks provide extensive support for custom dataset training, enabling the effective application of state-of-the-art detection models to our specialized dataset.
@@ 5.3 Benchmarking
Faster R-CNN [ren2015faster] 0.3656 0.6340 Dino [liu2023grounding] 0.4710 0.7890
RetinaNet [lin2017focal] 0.3834 0.6933 Rtmdet [lyu2022rtmdet] 0.4420 0.7080
Cascade R-CNN [cai2018cascade] 0.4860 0.7350 YOLOv5 [jocher2022ultralytics] 0.4920 0.8080
YOLOv9-e [wang2024yolov9] 0.6883 0.8927 YOLOv9-c [wang2024yolov9] 0.5562 0.7996
@@ 6.1 Model Optimization for Complex Disaster Scenarios
To investigate the impact of domain-specific training on model performance in complex disaster scenarios, we conducted a comparative analysis using several datasets and a model [nihal2024blurry]: C2A (synthetic disaster scenes), SARD (real-world search and rescue images), and "General Human Detection" (a combination of crowd human [shao2018crowdhuman], tiny person [yu2020scale], and VisDrone [du2019visdrone] datasets). By training models on these datasets and evaluating their performance across different validation sets, we aimed to identify the most effective approach for detecting humans in challenging disaster environments. Table 5 presents the results of this experiment, showcasing the mAP scores achieved by models trained on various datasets and validated on different test sets.
@@ 7 Conclusion
Our research presents the C2A dataset as a pivotal resource for enhancing machine learning models’ performance in UAV-assisted SAR operations within disaster-stricken environments. Despite its synthetic origins, the dataset serves as a critical step toward developing more accurate and robust detection systems. We have demonstrated that combining the C2A dataset with general human datasets leads to improved performance and generalization across various scenarios, highlighting the importance of integrating domain-specific and general data for effective SAR operations. The comparative analysis of model performance across different datasets underscores the value of the C2A dataset in enhancing the operational efficacy of search and rescue missions. However, we also acknowledge the limitations of relying solely on synthetic data and the need for incorporating real-world examples to increase confidence in the model’s real-world applicability. The limitations identified through our study provide a roadmap for future improvements, particularly in terms of scaling and positioning to mimic the unpredictable nature of real-world scenarios. We foresee the incorporation of actual disaster footage and advanced contextual understanding as key avenues for the next phase of dataset development. As the field progresses, the ultimate goal remains clear: to refine these tools and datasets to ensure that they can be relied upon when they are most needed—to save lives and navigate the complexities of disaster response with unparalleled precision and reliability.

### v2

@@ Abstract
Unmanned aerial vehicles (UAVs) have revolutionized search and rescue (SAR) operations, but the lack of specialized human detection datasets for training machine learning models poses a significant challenge. To address this gap, this paper introduces the Combination to Application (C2A) dataset, synthesized by overlaying human poses onto UAV-captured disaster scenes. Through extensive experimentation with state-of-the-art detection models, we demonstrate that models fine-tuned on the C2A dataset exhibit substantial performance improvements compared to those pre-trained on generic aerial datasets. Furthermore, we highlight the importance of combining the C2A dataset with general human datasets to achieve optimal performance and generalization across various scenarios. This points out the crucial need for a tailored dataset to enhance the effectiveness of SAR operations. Our contributions also include developing dataset creation pipeline and integrating diverse human poses and disaster scenes information to assess the severity of disaster scenarios. Our findings advocate for future developments, to ensure that SAR operations benefit from the most realistic and effective AI-assisted interventions possible. The dataset, code, and model are publicly available at: https://github.com/Ragib-Amin-Nihal/C2A
Keywords:
Aerial Object Detection UAV (Unmanned Aerial Vehicle) Human Detection Disaster Response Search and Rescue (SAR) Artificial Intelligence in Disaster Relief Emergency Management Benchmark Dataset
@@ 1 Introduction
The advancement of UAVs, colloquially known as drones, has signaled a new era in the field of emergency response and disaster management. With their unparalleled agility and ability to provide an aerial perspective, drones have rapidly become indispensable assets in the arsenal of SAR operations worldwide. These technological marvels significantly improve the efficiency and effectiveness of missions aimed at locating and aiding people in disaster-hit areas [13]. Drones can have a significant impact on minimizing the aftermath of disasters through time efficiency, making a difference in survival and fatality rates.
@@
In critical scenarios, the lack of datasets with the necessary detail for training advanced machine learning algorithms hinders the optimization of drone capabilities. Recognizing the critical need for a specialized dataset, our research introduces a novel dataset explicitly designed to improve human detection capabilities in disaster scenarios through drone vision. This dataset is created by combining human posture images with disaster scene backgrounds, resulting in an intricate collection of images that simulate the diverse settings faced during real-world SAR operations. By integrating human figures into a variety of disaster scenes, our approach aims to mirror the array of challenges that SAR drones are likely to face, facilitating the development of more robust and effective machine learning models tailored for disaster response applications [21].
Our endeavor is inspired by the growing field of disaster management technology, where recent breakthroughs highlight the revolutionary potential of machine learning and computer vision in improving SAR operations [4]. Our research focuses on creating a challenging dataset to train AI models for detecting partially occluded humans, a common scenario in disaster-stricken environments. Such occlusions, resulting from individuals being trapped under debris or obscured by various objects, represent a significant hurdle for human detection systems. Existing datasets scarcely address this challenge, primarily due to the inherent difficulty in replicating these complex scenarios accurately. Our dataset, therefore, signifies a pioneering step towards addressing this gap, offering a resource that simulates the conditions under which SAR operations unfold in reality.
We aim to enhance the precision and reliability of drone-operated human detection in emergency scenarios, underscoring the nuanced requirements of effective disaster response strategies. The construction of this dataset involved a process of incorporating elements of obstruction to emulate the visibility challenges frequently encountered in real-world disaster settings. This approach facilitates the training of AI models that are not only adept at recognizing human forms in clear view, but are also capable of inferring the presence of individuals in less than optimal visibility conditions. Such abilities are necessary for quickly finding emergency areas and finding survivors who need help right away, which makes SAR missions much more effective [2, 8].
@@ 2 Literature Review
Robust datasets are fundamental to developing and training precise machine learning models. The importance of datasets tailored to specific application domains is well established in the literature. Studies such as those by [1] and [27] emphasize the need for datasets that encompass the intricacies of various disaster scenarios. These resources are pivotal for calibrating UAV-operated detection systems to recognize human subjects under a multitude of conditions. Yet, there remains a scarcity of datasets that accurately mirror the complexities of disaster-hit environments.
Debris and destruction frequently obscure human subjects in disaster zones, biasing UAV detection. According to the literature, existing datasets and models are progressing, but they fall short of providing the granular detail required for reliable detection in such complex circumstances [19, 18]. A significant body of research, including work by [11], emphasizes the pressing need for advancing UAV technology to navigate these obstacles adeptly. Yet, the development of datasets that reflect the reality of partial occlusions in disaster contexts is still in its early stages, indicating a pivotal area for future research.
The evolution of machine learning and computer vision has been enhancing to UAV capabilities. These advancements have paved the way for more nuanced data analysis, crucial for discerning human presence within complex terrains. Even with these improvements, the research shows that algorithms and models still need to be improved, especially to make human detection more reliable [12].
One notable effort in this direction is the Search and Rescue Drone (SARD) dataset [25], which focuses on human detection in search and rescue operations using drone imagery. The SARD dataset includes images of people in various poses simulating exhausted or injured individuals, captured in non-urban environments. While it provides a valuable resource for developing detection models for SAR scenarios, the SARD dataset does not explicitly include disaster scenes, which pose additional challenges such as debris, occlusions, and clutter. Moreover, although the dataset contains diverse human poses, this information is not directly incorporated into the annotations, limiting its utility for pose-aware detection.
@@
Feature C2A (proposed) LoveNAS [30] Smoke Scene [31] M4SFWD [29] SARD [25]
@@ 3.1 Data Sources and Composition
AIDER (Aerial Image Dataset for Emergency Response Applications): The AIDER dataset [14] serves as the foundation for the disaster scene backgrounds. It comprises images from four major disaster types: Fire/Smoke (320 images), Flood (370 images), Collapsed Building/Rubble (320 images), and Traffic Accidents (335 images). These authentic disaster images offer a realistic portrayal of the chaotic and unpredictable conditions typical in emergency scenarios. We did not utilize the 1,200 normal case images to keep the focus on emergency situations. This dataset offers a glimpse into the chaotic and unpredictable environments that characterize disaster scenes, making it an ideal choice for our purposes.
LSP/MPII-MPHB (Multiple Poses Human Body): For the human subjects, we sourced images from the LSP/MPII-MPHB dataset [3, 10], which contains 26,675 images featuring 29,732 instances of human bodies in various poses. This dataset is specifically designed to capture a wide range of human body positions, including bent, kneeling, sitting, upright, and lying, providing the necessary diversity to train models for detecting humans under different conditions. The detailed annotations of human poses in this dataset are critical for training models to recognize human figures in complex disaster environments.
@@ 3.2 Pipeline Steps
Using the U2Net segmentation model [23], we isolated human figures from the LSP/MPII-MPHB dataset by removing the background. The U2-Net, short for "U-Squared Net," is a deep neural network known for its powerful performance in salient object detection and image segmentation tasks. It employs a nested U-structure that enhances the learning of local and global features within images, enabling precise segmentation of objects, including human figures, from their backgrounds. This process involved saving each figure with its respective pose in a separate folder, ensuring that the focus remained on the human subject without any background distractions.
@@ 4 Properties of C2A Dataset
The C2A (Combination to Application) dataset [^1] Dataset available at: https://github.com/Ragib-Amin-Nihal/C2A is a curated collection specifically designed for advancing human detection disaster scenarios by combining AIDER dataset images (disaster scene backgrounds) and diverse human poses from the LSP/MPII-MPHB dataset. Some of the samples of the dataset are presented in Fig. 1. Comparison of various datasets is shown on Table 2. In this section, we present a comprehensive analysis of the dataset’s properties.
Table 2: Comparison of various datasets [33] including the proposed C2A dataset
@@ 4.2 Objects Size
Table 3: Comparison of instance size (in terms of width) distribution of some datasets in aerial images and natural images. Some statistics collected from [33]
@@ 5.2 Training Options
The evaluation of the models on the C2A dataset was conducted using NVIDIA A100 GPUs, with a uniform batch size of 24 and an image resolution of 640x640 pixels across 50 epochs. The ADAM optimizer was chosen for its efficiency in handling large datasets and complex image structures. Basic data augmentation techniques, such as flipping and resizing, were employed to enhance model robustness and prevent overfitting. The experiments were facilitated by popular deep learning frameworks, specifically mmDetection [6], Detectron2 [32], and Ultralytics [9], known for their high performance in object detection tasks. These frameworks provide extensive support for custom dataset training, enabling the effective application of state-of-the-art detection models to our specialized dataset.
@@ 5.3 Benchmarking
Faster R-CNN [24] 0.3656 0.6340 Dino [16] 0.4710 0.7890
RetinaNet [15] 0.3834 0.6933 Rtmdet [17] 0.4420 0.7080
Cascade R-CNN [5] 0.4860 0.7350 YOLOv5 [9] 0.4920 0.8080
YOLOv9-e [28] 0.6883 0.8927 YOLOv9-c [28] 0.5562 0.7996
@@ 6.1 Model Optimization for Complex Disaster Scenarios
To investigate the impact of domain-specific training on model performance in complex disaster scenarios, we conducted a comparative analysis using several datasets and a model [20]: C2A (synthetic disaster scenes), SARD (real-world search and rescue images), and "General Human Detection" (a combination of crowd human [26], tiny person [34], and VisDrone [7] datasets). By training models on these datasets and evaluating their performance across different validation sets, we aimed to identify the most effective approach for detecting humans in challenging disaster environments. Table 5 presents the results of this experiment, showcasing the mAP scores achieved by models trained on various datasets and validated on different test sets.
@@ 7 Conclusion
In the rapidly changing field of disaster response, our research introduces the C2A dataset as a crucial resource, connecting AI with humanitarian efforts. This work advances the technical capabilities of UAV-assisted search and rescue operations and represents a significant shift in how we integrate machine learning into crisis management.
The C2A dataset fills a critical need in disaster response, offering a comprehensive, synthetic environment that represents the complexities of real-world catastrophes. This dataset forms a foundation for training more robust and adaptable AI models. Our comparative analysis across various datasets highlights the importance of combining domain-specific data (like C2A) with general human datasets, resulting in models that are both specialized and widely applicable.
These advancements are initial steps in an ongoing process. As we expand the capabilities of AI-assisted disaster response, we must consider the ethical implications and real-world applicability of our work. The limitations we’ve identified, particularly in synthetic data generation and real-world validation, serve as guides for future research.
We envision a future where AI becomes an essential tool in crisis management, working alongside human expertise to save lives and reduce suffering. To achieve this vision, we encourage the research community to:
  1. 1.
    Test the C2A dataset in real-world pilot studies, linking synthetic training with practical application.
  2. 2.
    Collaborate to expand and refine the dataset, including diverse disaster scenarios and cultural contexts.
  3. 3.
    Pursue research that combines computer vision, disaster management, and ethics to ensure responsible and effective use of AI in humanitarian efforts.
As the field advances, our ultimate goal goes beyond technological progress; we aim to create tools and methods that are reliable in critical situations. By continually improving our approach, incorporating real-world feedback, and fostering collaboration across disciplines, we move towards a future where technology and human compassion work together, addressing the complexities of disaster response with increased precision and reliability.
In conclusion, the C2A dataset and our findings represent not only a technical achievement but also progress towards a more resilient and responsive global community. As we confront increasingly complex global challenges, the combination of AI and human ingenuity offers hope for more effective, efficient, and compassionate disaster response strategies in the years to come.
Acknowledgements
This work was supported by JSPS KAKENHI Grant No. JP22F22769 and JP22KF0141. Also, this work was performed the commissioned research fund provided by F-REI (JPFR23010102).